Carmen is 6 years older than mike. In two years Carmen will be twice as old as mike. What are their ages?
Answer by mananth(16946)

let mike be x years
carmen will be x+6
..
ages after 2 years
mike = x+2
carmen x+6+2
..
equation after 2 years
x+8=2(x+2)
x+8=2x+4
2x-x=8-4
x=4 , mike's age
carmen - 10 years